# Spectra-to-Structure and Structure-to-Spectra Inference Across the Periodic Table

### v1

@@ A.1 XAFS Theoretical Background
Figure S1 presents a typical Cu K-edge X-ray absorption spectrum, illustrating the fundamental decomposition of the total absorption signal into the near-edge (XANES) and extended (EXAFS) regions. The sharp increase in absorption around 8980-8990 eV corresponds to the ionization of a Cu 1s core electron and marks the absorption edge. The edge jump $\Delta\mu_{\text{Cu}}$ quantifies the magnitude of this transition and is used for normalization in EXAFS analysis.
To the left of the edge, a weak but resolvable pre-edge feature appears, often associated with 1s $\rightarrow$ 3d or 4p transitions facilitated by hybridization or local symmetry breaking. The red inset illustrates this as a local Cu-O cluster. To the right of the edge, the XANES region spans roughly 30-50 eV and encodes multiple-scattering, oxidation state, and coordination effects. The EXAFS region follows, showing oscillations in $\mu(E)$ caused by interference between the outgoing photoelectron wave and waves backscattered by neighboring atoms. These oscillations decay with energy and distance due to the finite photoelectron mean free path.
@@ A.3 Experiment Results for XANES/EXAFS Predictions
These results underscore the benefit of XAStruct’s architecture in capturing nuanced structure-spectrum relationships, thanks to its domain-informed input encoding and unified learning framework. Nonetheless, the variations in prediction accuracy across elements also suggest space for further optimization, particularly in handling underrepresented edge types or elements with sparse training data.
@@ A.4 Experiment Results for Structure Descriptor Predictions
Figures S6 and S7 evaluate the performance of XAStruct and a baseline RF model in predicting MNND from XAFS spectra from the edge of the K-edge. Figure S6 presents the mean absolute errors (MAEs) of elements on the $\text{log}_{10}$ scale, while Figure S7 shows a correlation scatter plot between the predicted and true MNND values in the dataset.
@@
[…]
@@
(a) Comparison of Accuracies
(b) Comparison of F-1 scores
Figure S8: Element-wise Performance Comparison for CN Prediction.
@@
Figure S8 presents a radar plot comparing the element-wise prediction performance of two CN classification models: a lightweight random forest (RF, used in XAStruct) and a baseline multi-layer perceptron (MLP). Figure 8(a) shows the accuracy per element, while Figure 8(b) shows corresponding F-1 scores.
@@
(a) Comparison of Accuracies
(b) Comparison of F-1 scores
Figure S9: Element-wise Performance Comparison for Neighbor Atom Prediction.
@@
Figure S9 compares the performance of our XAStruct model and an RF baseline in predicting the identity of neighboring atoms from XAS spectra. Figure 9(a) shows element-wise prediction accuracies, while Figure 9(b) displays F-1 scores. This task is framed as a multi-class classification problem where each absorbing atom is assigned a most probable neighbor type.
@@
However, some limitations remain for the neighbor atom prediction model. As visible in Figure 9(b), some elements, particularly those with broad coordination environments or low data representation, still show moderate to low F-1 scores. Additionally, while XAStruct predicts dominant neighbor types with high confidence, it does not capture the full neighbor distribution or handle multi-species environments. Future work may benefit from more expressive modeling of neighbor atom uncertainty, attention-based interaction learning, or probabilistic multi-label outputs to handle complex chemical environments.

### v2

@@ A.1 XAFS Theoretical Background
Figure S1 presents a typical Cu K-edge X-ray absorption spectrum, illustrating the fundamental decomposition of the total absorption signal into the near-edge (XANES) and extended (EXAFS) regions. The sharp increase in absorption around 8980–8990 eV corresponds to the ionization of a Cu 1s core electron and marks the absorption edge. The edge jump $\Delta\mu_{\text{Cu}}$ quantifies the magnitude of this transition and is used for normalization in EXAFS analysis.
To the left of the edge, a weak but resolvable pre-edge feature appears, often associated with 1s $\rightarrow$ 3d or 4p transitions facilitated by hybridization or local symmetry breaking. The red inset illustrates this as a local Cu–O cluster. To the right of the edge, the XANES region spans roughly 30–50 eV and encodes multiple-scattering, oxidation state, and coordination effects. The EXAFS region follows, showing oscillations in $\mu(E)$ caused by interference between the outgoing photoelectron wave and waves backscattered by neighboring atoms. These oscillations decay with energy and distance due to the finite photoelectron mean free path.
@@ A.3 Experiment Results for XANES/EXAFS Predictions
These results underscore the benefit of XAStruct’s architecture in capturing nuanced structure-spectrum relationships, thanks to its domain-informed input encoding and comprehensive learning framework. Nonetheless, the variations in prediction accuracy across elements also suggest space for further optimization, particularly in handling underrepresented edge types or elements with sparse training data.
@@ A.4 Experiment Results for Structure Descriptor Predictions
Figures S6 and S7 evaluate the performance of XAStruct and a baseline RF model in predicting MNND from K-edge XAFS spectra. Figure S6 presents the mean absolute errors (MAEs) of elements on the $\text{log}_{10}$ scale, while Figure S7 shows a correlation scatter plot between the predicted and true MNND values in the dataset.
@@
Figure S7 further highlights the regression behavior of both models. XAStruct produces a tight distribution around the identity line with an $R^{2}$ of $0.985$ and MAE of $0.0350\mathring${A}. In contrast, the RF baseline shows increased scatter and a slightly lower $R^{2}$ of $0.983$ and much larger MAE of $0.0557\mathring${A}. As visualized in the figure and quantified in our main text, the RF model exhibits a consistent bias: it underestimates large MNND values and overestimates smaller ones, resulting in a regression slope of $0.92$. XAStruct, by contrast, achieves a slope of $0.97$, indicating improved fidelity in modeling the true linear relationship.
@@
Figure S8: Element-wise Performance Comparison for CN Prediction Top-1 Accuracies
Figure S9: Element-wise Performance Comparison for CN Prediction F-1 Scores.
@@
Figure S8, S9 present radar plots comparing the element-wise prediction performance of two CN classification models: a lightweight random forest (RF, which is a component of the XAStruct framework) and a baseline multi-layer perceptron (MLP). Figure S8 shows the accuracy per element, while Figure S9 shows corresponding F-1 scores.
@@
Figure S10, S11 compares the performance of the deep learning component of our XAStruct framework and an RF baseline in predicting the identity of neighboring atoms from XAS spectra. Figure S10 shows element-wise prediction accuracies, while Figure S11 displays F-1 scores. This task is framed as a multi-class classification problem where each absorbing atom is assigned a most probable neighbor type.
@@
However, some limitations remain for the neighbor atom prediction model. As visible in Figure S11, some elements, particularly those with broad coordination environments or low data representation, still show moderate to low F-1 scores. Additionally, while XAStruct predicts dominant neighbor types with high confidence, it does not capture the full neighbor distribution or handle multi-species environments. Future work may benefit from more expressive modeling of neighbor atom uncertainty, attention-based interaction learning, or probabilistic multi-label outputs to handle complex chemical environments.
Figure S10: Element-wise Performance Comparison for Neighbor Atom Prediction Top-1 Accuracies.
Figure S11: Element-wise Performance Comparison for Neighbor Atom Prediction F-1 Scores.